# Language Models for Controllable DNA Sequence Design

### v1

@@ 5.2 Enhancer Generation
Dataset and task descriptions. This task involves two distinct datasets: one from fly brain (Janssens et al., 2022) and the other from human melanoma cells (Atak et al., 2021). We follow the same data split as in the work by Stark et al. (2024), using 104k fly brain DNA sequences and 89k human melanoma DNA sequences, each consisting of 500 base pairs. The cell class labels are derived from the ATAC-seq experiments (Buenrostro et al., 2013), with 47 classes for the human melanoma dataset and 81 classes for the fly brain dataset. Our objective is to generate DNA sequences based on the given cell class labels.
Metrics. Following Stark et al. (2024), we evaluate the performance using the Fréchet Biological Distance (FBD) between the generated and original DNA sequences. To calculate the FBD score, we use a pretrained classifier model (Stark et al., 2024) to obtain the hidden representations and then compute the Wasserstein distance between Gaussians. Additionally, we assess performance using the diversity score and fluency score, similar to the metrics proposed in Section 4. These metrics provide a comprehensive evaluation of the generated DNA sequences.
@@ 5.3 ChIP-Seq Generation
6 Summary and Discussion
In this paper, we present ATGC-Gen, a controllable DNA sequence generator that integrates diverse biological properties via cross-modal encoding and leverages language models for conditional generation. Our method enables accurate, fluent, and biologically relevant sequence design across multiple tasks. To support this direction, we also construct a new dataset from ChIP-Seq experiments, providing a realistic benchmark for DNA-protein binding generation.
@@ 6 Summary and Discussion
Broader Impacts. Controllable DNA generation offers exciting opportunities for synthetic biology and genetic design. While it may accelerate the development of beneficial traits, it also raises concerns about misuse or the unintended synthesis of harmful sequences. Ensuring biological safety and ethical safeguards is critical.
Future Directions. Future work may explore more advanced generation techniques (e.g., retrieval-augmented models), as well as the development of richer, biologically grounded benchmarks to drive progress in programmable genomics.
@@ Appendix A Data Processing for ChIP-Seq Experiment
Embedding Protein Sequences. To represent the transcription factor (TF) proteins, we use the ESM-2-3B protein language model (Lin et al., 2023) to extract contextual embeddings from the amino acid sequences. We retain the full-length output embeddings without applying any sequence-level pooling or aggregation, thereby preserving residue-level resolution for downstream tasks.
@@ C.1 Enhancer Design without Property Information
To better understand the role of conditioning signals, we follow prior work and evaluate ATGC-Gen-GPT in an uncontrolled setting, where no property information is provided during generation (Stark et al., 2024). As shown in Table 6, ATGC-Gen-GPT results in a mixed performances compared with the Dirichlet FM baseline in this scenario. This result contrasts with the controllable generation setting, where ATGC-Gen-GPT consistently outperforms Dirichlet FM by a large margin (Table 3). The discrepancy highlights that ATGC-Gen is particularly effective when it can leverage external biological properties to guide sequence generation. In the absence of such conditioning, its advantage diminishes. Overall, these findings suggest that while Dirichlet FM remains competitive in unconditional generation, ATGC-Gen is better suited for property-conditioned tasks—precisely the setting that is most relevant in practical genomic design applications.

### v2

@@ 5.2 Enhancer Generation
Dataset and task descriptions. This task involves two distinct datasets: one from fly brain (flybrain) and the other from human melanoma cells (melanoma). We follow the same data split as in the work by stark2024dirichlet, using 104k fly brain DNA sequences and 89k human melanoma DNA sequences, each consisting of 500 base pairs. The cell class labels are derived from the ATAC-seq experiments (buenrostro2013transposition), with 47 classes for the human melanoma dataset and 81 classes for the fly brain dataset. Our objective is to generate DNA sequences based on the given cell class labels.
Metrics. Following stark2024dirichlet, we evaluate the performance using the Fréchet Biological Distance (FBD) between the generated and original DNA sequences. To calculate the FBD score, we use a pretrained classifier model (stark2024dirichlet) to obtain the hidden representations and then compute the Wasserstein distance between Gaussians. Additionally, we assess performance using the diversity score and fluency score, similar to the metrics proposed in Section 4. These metrics provide a comprehensive evaluation of the generated DNA sequences.
@@ 5.3 ChIP-Seq Generation
6 Conclusion and Discussion
In this paper, we present ATGC-Gen, a controllable DNA sequence generator that integrates diverse biological properties via cross-modal encoding and leverages language models for conditional generation. Our method enables accurate, fluent, and biologically relevant sequence design across multiple tasks. To support this direction, we also construct a new dataset from ChIP-Seq experiments, providing a realistic benchmark for DNA–protein binding generation.
Practical Usage and Applications. Controllable sequence design is closely aligned with real-world needs in synthetic biology, such as promoter or enhancer engineering and transcription-factor binding site optimization. ATGC-Gen can assist researchers in generating candidate sequences that satisfy multiple design constraints before downstream experimental screening. Although this work does not include wet-lab validation, the proposed controllability interface aligns with how sequence constraints are applied in practical design settings.
@@ 6 Conclusion and Discussion
Broader Impacts. Controllable DNA generation offers exciting opportunities for synthetic biology and genetic engineering. While such capabilities may accelerate the design of beneficial biological functions, they also raise concerns about misuse or the unintended synthesis of harmful sequences. Ensuring biological safety, responsible use, and ethical safeguards remains critical.
Future Directions. Future work may explore more advanced generation techniques such as retrieval-augmented models, as well as the development of richer and biologically grounded benchmarks to drive progress in programmable genomics. Incorporating experimental feedback loops or active-learning strategies may further enhance the practical utility of controllable DNA generation.
@@ Appendix A Data Processing for ChIP-Seq Experiment
Embedding Protein Sequences. To represent the transcription factor (TF) proteins, we use the ESM-2-3B protein language model (lin2023evolutionary) to extract contextual embeddings from the amino acid sequences. We retain the full-length output embeddings without applying any sequence-level pooling or aggregation, thereby preserving residue-level resolution for downstream tasks.
@@ C.1 Enhancer Design without Property Information
To better understand the role of conditioning signals, we follow prior work and evaluate ATGC-Gen-GPT in an uncontrolled setting, where no property information is provided during generation (stark2024dirichlet). As shown in Table 6, ATGC-Gen-GPT results in a mixed performances compared with the Dirichlet FM baseline in this scenario. This result contrasts with the controllable generation setting, where ATGC-Gen-GPT consistently outperforms Dirichlet FM by a large margin (Table 3). The discrepancy highlights that ATGC-Gen is particularly effective when it can leverage external biological properties to guide sequence generation. In the absence of such conditioning, its advantage diminishes. Overall, these findings suggest that while Dirichlet FM remains competitive in unconditional generation, ATGC-Gen is better suited for property-conditioned tasks—precisely the setting that is most relevant in practical genomic design applications.
@@ C.3 Effect of Unmasking Rate on ATGC-Gen-Bert
Figure 4: Fraction of generated sequences containing at least one occurrence of each selected transcription factor motif. Highly enriched motifs include GABPA, KLF4, and YY1.
Figure 5: Motif distribution across five randomly sampled generated sequences. Each box marks a detected motif occurrence at the corresponding genomic position.
C.4 Motif Enrichment in Generated Sequences
To further validate the biological plausibility of our generated regulatory sequences, we performed motif enrichment analysis using a curated panel of transcription factor (TF) binding motifs from the JASPAR database (rauluseviciute2024jaspar) on the promoter design task described in Section 5.1. Specifically, we considered ten motifs corresponding to well-established regulators of gene expression:
  • SP1 (MA0079.5)
  • KLF4 (MA0039.5)
  • NRF1 (MA0506.3)
  • YY1 (MA0095.4)
  • NFYA (MA0060.4)
  • ELK1 (MA0028.3)
  • ETS1 (MA0098.4)
  • GABPA (MA0062.4)
  • CREB1 (MA0018.5)
  • E2F1 (MA0024.3)
We scanned all generated sequences for the presence of these motifs using position frequency matrices (PFMs) from JASPAR, followed by enrichment quantification.
Global motif enrichment.
Figure 4 summarizes the fraction of generated sequences that contain at least one occurrence of each TF motif. Highly enriched motifs include GABPA, KLF4, and YY1, with more than $80\%$ of sequences exhibiting at least one instance. This demonstrates that the model captures essential cis-regulatory elements commonly associated with enhancer-like activity.
Motif positional distribution.
To illustrate the positional organization of motifs, Figure 5 shows motif annotations across five randomly selected generated sequences. Each motif occurrence is plotted at its genomic position, highlighting the combinatorial nature of motif clustering. The recurrence of canonical enhancer-associated motifs such as ETS1, SP1, and NFYA suggests that the generative model produces realistic cis-regulatory grammar beyond single-motif enrichment.